# Hypersonic flow onto a large curved wedge and the dissipation of shock wave

### v1

@@ Abstract.
For supersonic flow past an obstacle, experiments show that the flow field after shocks changes slightly for incoming flow with Mach number larger to 5, named hypersonic flow. Hypersonic similarity principle was first found by Qian for thin wedge by studying potential flow. In this paper, we explore the existence of smooth flow field after shock for hypersonic potential flow past a curved smooth wedge with neither smallness assumption on the height of the wedge nor that it is a BV perturbation of a line. The asymptotic behaviour of the shock is also analysed. We proved that for given Bernoulli constant of the incoming flow, there exists a sufficient large constant such that if the Mach number of the incoming flow is larger than it, then there exists a global shock wave attached to the tip of the wedge together with a smooth flow field between it and the wedge. The state of the flow after shock is in a neighbourhood of a curve that is determined by the wedge and the density of the incoming flow. If the slope of the wedge has a positive limit as $x$ goes to infinity, then the slope of the shock tends to that of the self-similar case that the same incoming flow past a straight wedge with slope of the limit. Specifically, we demonstrate that if the slope of the wedge is parallel to the incoming flow at infinity, then the strength of the shock will attenuate to zero at infinity. The restrictions on the surface of a wedge have been greatly relaxed compared to the previous works on supersonic flow past wedges. The method employed in this paper is characteristic decomposition, and the existence of the solution is obtained by finding an invariant domain of the solution based on geometry structures of the governing equations. The idea and the method used here may be helpful for other problems.
@@
hypersonic flow; piecewise smooth potential flow; global attached shock; shock wave dissipation;
@@ 1. Introduction
We are concerned with the problems of attached shock arising in hypersonic flows past wedges with large data. These problems are fundamental not only in aerodynamics but also in the mathematical theory of multidimensional conservation laws since their solutions plays the role of building blocks of general solutions to the multidimensional Euler equations for compressible fluids [1] [27]. When a supersonic flow past a wedge with angle less than a critical value, attached shock appears ahead of it. As the Mach number of the upcoming flow increases, the shock fronts approach the wedge, the region between the shock and the wedge becomes narrower and narrower, which is called a thin shock layer. The flow is called hypersonic if its Mach number is bigger than 5. The hypersonic flows show difference features from that of supersonic flows. The thin shock layer experiences small variation as the Mach number of the incoming flow varies after it is larger than some value. This important feature was observed in physical and engineer. However, many fundamental issues for hypersonic flow past obstacles have not been understood, including the global structure, stability and asymptotic. Therefore, it is essential to establish the global existence and structural stability of solutions of hypersonic flow past an obstacle in order to understand fully the Mach independent principle. On the other hand, there has been no rigorous mathematical results on the global existence and stability of the solution, including the case of potential flow which is widely used in aerodynamics (cf. [1], [27]). One of the main reasons is that the problems involve several challenging difficulties in the analysis of nonlinear partial differential equations such as solution with large data, free boundary problems, and singularity similar as vacuum when the shock layer is thin. In this paper we develop a rigorous mathematical approach to overcome these difficulties and establish a global theory of existence and stability for hypersonic flow past a curved wedge for potential flow. The techniques and ideas developed here will be useful for other nonlinear problems involving similar difficulties.
On the $(x,y)$ plane, the steady Euler equations for potential flow are
@@
where the first line is the conservation law of mass and the second is the irrotational condition. The unknowns $\rho$ and $(u,v)$ denote respectively the density of the mass and the velocity of the flow. We use the notation $\vec{U}:=(u,v)$. For polytropic gas, the pressure of the flow is
@@
with $q^{2}=\sqrt{u^{2}+v^{2}}$. For (1.0.3), it can be solved that
@@
In the upper $(x,y)$ plane, there is a wedge W with a given $C^{2}$ smooth boundary function $y=f(x)$, where $f(x)>0$ for $x>0$ and $f(0)=0$. Here, we are mainly concerned with the following two kinds of wedge respectively.
@@
It is worth to note that the assumptions on $f$ in Case A means no small BV bounded or even any convexity.
@@
Our problem can be formulate as follows: Problem 1 (Initial boundary value problem): Given $\mathbf{B}$, seek a solution of system (1.0.1) with given velocity $\left(\underline{u},\underline{v}\right)$ of the incoming flow at $x=0$
@@
where $y=\phi(x)$ is the discontinuity of the solution, namely, the shock front, along which it holds the Rankine-Hugoniot condition
@@
Problem 2 (Free boundary value problem): Determine the flow fields $(u,v)$ and its asymptotic behaviour in the domain $\Omega$ bounded by the wedge W and the free boundary $\textsf{S}:y=\phi(x)$, satisfying (1.0.1), (1.0.3), (1.0.14), (1.0.15) and (1.0.16) for a given smooth function $f$ satisfying either Case A ((1.0.6) (1.0.7)) or Case B ((1.0.9) (1.0.10)). Meanwhile the entropy condition $p>\underline{p}$ should hold along the shock.
@@
We develop a rigorous mathematical approach to extend the local theory to a global theory for solutions of an attached shock. The shock approach to the wedge as the Mach number of the incoming flow $\underline{\mathbf{M}}$ tends to infinity. The solution is supersonic and $C^{1}$ smooth in the domain bounded by the shock and the wedge. The shock fronts is $C^{2}$ smooth. Our main results are stated as follows
@@
For the supersonic incoming flow with Mach number $\underline{\mathbf{M}}$ and fixed Bernoulli’s constant $\mathbf{B}$, there exists $\mathbf{M}_{1}>1$, such that, for $\underline{\mathbf{M}}>\mathbf{M}_{1}$, the problem (1.0.1), (1.0.13) and (1.0.14) admits a global piecewise smooth solution $(\rho,u,v)$ with an attached shock $y=\phi(x)$ satisfying (1.0.16), provided that the wedge boundary $f\in C^{2}(\mathbb{R}^{+})$ satisfies (1.0.6)-(1.0.8).
@@
where $(u_{\infty},v_{\infty})$ is the constant supersonic solution for corresponding problem the straight wedge $f(x)=f_{\infty}x$, and
@@
Furthermore, the shock will approaches to the wedge as the Mach number of the incoming flow increases.
@@
There are many literatures on the global existence of the entropy solutions of small data in the BV space for the one dimensional hyperbolic conservation laws since J. Glimm’s original paper [32] in 1960s. There are also many literatures on the BV solutions of the two dimensional steady supersonic Euler flow with small data. Due to the nonlinearity of the system, few results on the global existence of weak solutions with large data are obtained for system with special structures, see Nishida [58] for one dimensional isothermal gas dynamic system, Nishida-Smoller for the isentropic gas with $\gamma$ sufficiently close to 1 [59]).
Our proof originates from a fundamental idea. Under the assumption of an invariant Bernoulli constant, as the incoming flow continuously accelerates, approaching the limit speed, the corresponding density of the gas will decay to zero. Consequently, the associated mass flux diminishes as well. In this context, we can consider an asymptotic state defined as a limit solution, where the shock front coincides with the solid wall. The flow field after the shock is defined through a combination of the Rankine-Hugoniot conditions and the solid wall boundary conditions (see Figure 4). Then, our estimates rely on three fundamental principles. First, owing to the conservation of mass and the entropy condition, we can control the distance between the shock front and the solid wall. This provides us with an estimate of the shock (see Theorem 4.2). Second, under the assumption that the derivatives of the flow field are uniformly bounded, as the shock front approaches the solid wall, the flow field’s velocity can be approximated by the velocities obtained through the combination of the shock conditions and the solid wall conditions. This approximation allows for an estimate of the flow field (see Theorem 4.6). Third, as we approach the limit solution, an analysis of the characteristics’ reflection provides us with an estimate for the derivatives of the solution (see Theorem 4.9). We point out here that the first two principles are directly reflected in the definition of limit solution, which is the logic of conservation laws and has not appeared in previous proofs. The third logic stems from the equation.
@@
We need to analyse a solution with large variations, in the presence of both free and fixed boundaries. Due to the complexity of the objects we are calculating, the presentation of our calculations is crucial. Our approach relies on two key techniques. The first is characteristic decomposition. In reference [48], the authors introduce the technique of characteristic decomposition for quasi-linear hyperbolic systems, clearly demonstrating the differential relationships among various quantities of the system, and further constructing global solutions for rarefaction wave interactions that occur in high-dimensional Riemann problems, see [24, 35, 42, 43, 44, 46, 47, 49, 50, 52]. We employ their technique to investigate the equations and their corresponding boundary relations. The second technique is used to discuss the phenomenon of shock dissipation. As the incoming and post-shock flows near the limit speeds, due to the significant singularity of various quantities at this time, the calculation is difficult. In this part, we only consider the case where the adiabatic constant is $\gamma=2$. There are two benefits to doing this. First, in this case, the state equation, Bernoulli’s equation, and the Rankine-Hugoniot conditions have polynomial forms, which streamline the calculations. Second, it ensures that various quantities maintain good differentiability near the limit speed. Utilizing computational software under these conditions, we are well-equipped to derive the desired computation. We believe that the methods discussed herein can be further extended to yield more general results.
@@
The rest of this paper is organized as follows. In Section 2, we first decompose both the system and boundary conditions by making use of the characteristics and reformulate the problem into one in which $\rho$, $(u,v)$ and their derivatives are expressed by $c$ and the directive derivation along the characteristic, that is ¯±$\bar{\partial}^{\pm}c$. In Section 3, the limit solution is introduced for the case $\underline{u}=\bar{q}$. A particular curve was specified by analysing the jump conditions, the Rankine-Hugoniot, that the shock satisfies. Estimates are obtained here for the states curve $(u,v)$ that could be connected to a given hypersonic flow by a shock by analysing the continuity of a family of parametric curves with parameter relating to the Mach number of the incoming flow. In Section 4.1, we present some a priori estimates for the solution, whose states locate around the limit solution. The proof of Theorem 1.1 is given in Section 4. We establish the a priori estimates for the flow with boundary of Case A in this section and get the global existence of solution. The asymptotic behaviour of the solution is analysed in Section 4.3. Furthermore, we analyse the higher order derivatives of the solution and obtain their limit as the Mach number of the incoming flow goes to infinity in Section 4.4. Finally, for Case B, the proof of Theorem 1.2 is given in Section 5. In Section 5.1, we present the delicate states estimates around the shock polar. Then, under the convex assumption, we show the sign estimates of the solution and prove the global existence. The shock is proved extinct at infinity for $f_{\infty}=0$.
@@ 2.1. Preliminary
In the following discussion, sometimes for the sake of convenience, we list both $\vec{U}$ and $\epsilon$, but in fact, they share a constrain condition (2.1.2).
@@ 2.2. Characteristic decomposition
Next, we intend to derive the relation of ¯±$\bar{\partial}^{\pm}c$ on the shock $y=\phi(x)$ by (2.1.2). By (1.0.16), define the angle of shock as follows
@@
$g(\vec{U},\epsilon)$ is continuous with respect to both $\vec{U}$ and $\epsilon$ under the constrain (2.1.2). As before, we use the notation $g=g(\vec{U},\epsilon)$ in the sequel without confusion. Therefore, Problem 2 induces the derivatives relations (2.2.9), (2.3.2) and (2.4.8).
@@ 3. A limit case
In this section, we introduce the limit solution for the case where $\underline{u}=\bar{q}$ ($\underline{\rho}\underline{u}=\epsilon=0$) and present some useful estimates. This solution is first introduced in [33, 34] to consider the hypersonic flow pasting a bend obstacle.
@@ 3.1. The limit solution
On the $(u,v)$ plane, it is a circle pasting the origin and centring at $(\frac{\bar{q}}{2},0)$. For such a case, we have many simplification. First, for the shock $y=\phi(x)$, since $\underline{\rho}=\rho(\bar{q})=\bar{\rho}=0$, it yields
@@
We define a nonzero flow fields $(u_{s},v_{s})$ along $\mathsf{W}$ by combining (1.0.14) and (3.1.3) directly as follows
@@
where $S_{o}>0$ is any positive constant. Based on this limit solution, the authors establish global existence of detach transonic shock for a supersonic flow onto a blunt body in [3].
@@ 3.2. Differential relations for the limit solution
Moreover, the shock polar (2.1.2) becomes the circle (2.4.5), so we have
@@
We can calculate directly, for the limit solution $\left(u_{s},v_{s}\right)$ in (3.1.5), (3.2.4) and (3.2.6) are equivalent to
@@
If $0<\underline{f}<\bar{f}<\sqrt{\frac{2}{\gamma-1}}$, then there exist a positive constant $\delta_{\bar{K}}$, which is depending on $\underline{f}$ and $\bar{f}$, to ensure the following estimate
@@
provided $0<\underline{f}<\bar{f}<\sqrt{\frac{2}{\gamma-1}}$. For any $(u,v)\in\mathcal{C}_{s}^{\delta_{\bar{K}}}$, similar as the proof in Lemma 3.4, we can get the conclusion by the continuity of the function $g(\vec{U},\epsilon)$. ∎

### v2

@@ Abstract.
The flow field with a Mach number larger than 5 is named hypersonic flow. In this paper, we explore the existence of smooth flow field after shock for hypersonic potential flow past a curved smooth wedge with neither smallness assumption on the height of the wedge nor that it is a BV perturbation of a line. The asymptotic behaviour of the shock is also analysed. We prove that for given Bernoulli constant of the incoming flow, there exists a sufficient large constant such that if the Mach number of the incoming flow is larger than it, then there exists a global shock wave attached to the tip of the wedge together with a smooth flow field between it and the wedge. The state of the flow after shock is in a neighbourhood of a curve that is determined by the wedge and the density of the incoming flow. If the slope of the wedge has a positive limit as $x$ goes to infinity, then the slope of the shock tends to that of the self-similar case that the same incoming flow past a straight wedge with slope of the limit. Specifically, we demonstrate that if the slope of the wedge is parallel to the incoming flow at infinity, the strength of the shock will diminish to zero at infinity. The restrictions on the surface of a wedge have been greatly relaxed compared to the previous works on supersonic flow past wedges. The method employed in this paper is characteristic decomposition, and the existence of the solution is obtained by finding an invariant domain of the solution based on geometry structures of the governing equations. The ideas and methods presented here may be applicable to other problems.
@@
Hypersonic flow; piecewise smooth potential flow; global attached shock; shock wave dissipation
@@ 1. Introduction
We are concerned with the problems of attached shock arising in hypersonic flows past wedges with large data. These problems are fundamental not only in aerodynamics but also in the mathematical theory of multidimensional conservation laws since their solutions play the role of building blocks of general solutions to the multidimensional Euler equations for compressible fluids [1] [27]. When a supersonic flow passes a wedge with an angle less than a critical value, an attached shock appears ahead of it. As the Mach number of the upcoming flow increases, the shock fronts approach the wedge, the region between the shock and the wedge becomes narrower and narrower, which is called a thin shock layer. The flow is called hypersonic if its Mach number is bigger than 5. The hypersonic flows show different features from those of supersonic flows. The thin shock layer experiences small variations as the Mach number of the incoming flow increases beyond a certain value. This important feature was observed in physical and engineering. However, many fundamental issues for hypersonic flow past obstacles have not been understood, including the global structure, stability and asymptotic behavior. Therefore, it is essential to establish the global existence and structural stability of solutions of hypersonic flow past an obstacle in order to understand fully the Mach number independence principle. On the other hand, there has been no rigorous mathematical results on the global existence and stability of the solution, including the case of potential flow which is widely used in aerodynamics (cf. [1] [27]). One of the main reasons is that the problems involve several challenging difficulties in the analysis of nonlinear partial differential equations such as solutions with large data, free boundary problems, and singularity similar as vacuum when the shock layer is thin. In this paper we develop a rigorous mathematical approach to overcome these difficulties and establish a global theory of existence and stability for hypersonic flow past a curved wedge for potential flow. The techniques and ideas developed here will be useful for other nonlinear problems involving similar difficulties.
On the $(x,y)$-plane, the steady Euler equations for potential flow are
@@
where the first line is the conservation of mass and the second is the irrotational condition. The unknowns $\rho$ and $(u,v)$ denote respectively the density of the mass and the velocity of the flow. We use the notation $\vec{U}:=(u,v)$. For polytropic gas, the pressure of the flow is
@@
with $q=\sqrt{u^{2}+v^{2}}$. For (1.0.3), it can be solved that
@@
In the upper half of the $(x,y)$ plane, there is a wedge W with a given $C^{2}$ smooth boundary function $y=f(x)$, where $f(x)>0$ for $x>0$ and $f(0)=0$. Here, we are mainly concerned with the following two kinds of wedge respectively.
@@
It is worth noting that the assumptions on $f$ in Case A mean no small BV bounded or even any convexity.
@@
Our problem can be formulated as follows: Problem 1 (Initial boundary value problem): Given $\mathbf{B}$, seek a solution of system (1.0.1) with given velocity $\left(\underline{u},\underline{v}\right)$ of the incoming flow at $x=0$
@@
where $y=\phi(x)$ is the discontinuity of the solution, namely, the shock front, along which it holds the Rankine-Hugoniot (R-H) condition
@@
Problem 2 (Free boundary value problem): Determine the flow fields $(u,v)$ and their asymptotic behaviour in the domain $\Omega$ bounded by the wedge W and the free boundary $\textsf{S}:y=\phi(x)$, satisfying (1.0.1), (1.0.3), (1.0.14), (1.0.15) and (1.0.16) for a given smooth function $f$ satisfying either Case A ((1.0.6) (1.0.7)) or Case B ((1.0.9) (1.0.10)). Meanwhile, the entropy condition $p>\underline{p}$ should hold along the shock.
@@
We develop a rigorous mathematical approach to extend the local theory to a global theory for solutions of an attached shock. The shock approaches to the wedge as the Mach number of the incoming flow $\underline{\mathbf{M}}$ tends to infinity. The solution is supersonic and $C^{1}$ smooth in the domain bounded by the shock and the wedge. The shock fronts is $C^{2}$ smooth. Our main results are stated as follows
@@
For the supersonic incoming flow with fixed Bernoulli’s constant $\mathbf{B}$, there exists $\mathbf{M}_{1}>1$ only depending on the wall and $\mathbf{B}$, such that, for $\underline{\mathbf{M}}>\mathbf{M}_{1}$, the problem (1.0.1), (1.0.13) and (1.0.14) admits a global piecewise smooth solution $(\rho,u,v)$ with an attached shock $y=\phi(x)$ satisfying (1.0.16), provided that the wedge boundary $f\in C^{2}(\mathbb{R}^{+})$ satisfies (1.0.6)-(1.0.8).
@@
where $(u_{\infty},v_{\infty})$ is the constant supersonic solution for corresponding problem involving the straight wedge $f(x)=f_{\infty}x$. Additionally, we have
@@
Furthermore, the shock approaches the wedge as the Mach number of the incoming flow increases.
@@
Our proof originates from a fundamental idea. Under the assumption of an invariant Bernoulli constant, as the incoming flow continuously accelerates, approaching the limit speed, the corresponding density of the gas will decay to zero. Consequently, the associated mass flux diminishes as well. In this context, we can consider an asymptotic state defined as a limit solution, where the shock front coincides with the solid wall. The flow field after the shock is defined through a combination of the Rankine-Hugoniot conditions and the solid wall boundary conditions (see Figure 4). Then, our estimates rely on three fundamental steps. First, owing to the conservation of mass and the entropy condition, we can control the distance between the shock front and the solid wall. This provides us with an estimate of the shock (see Theorem 4.2). Second, under the assumption that the derivatives of the flow field are uniformly bounded, as the shock front approaches the solid wall, the flow field’s velocity can be approximated by the velocities obtained through the combination of the shock conditions and the solid wall conditions. This approximation allows for an estimate of the flow field (see Theorem 4.6). Third, as we approach the limit solution, an analysis of the characteristics’ reflection provides us with an estimate for the derivatives of the solution (see Theorem 4.9). We point out here that the first two steps are directly reflected in the definition of limit solution, which is the logic of conservation laws and has not appeared in previous proofs. The third logic stems from the equation.
@@
We need to analyse a solution with large variations, in the presence of both free and fixed boundaries. Due to the complexity of the objects we are calculating, the presentation of our calculations is crucial. Our approach relies on two key techniques. The first is characteristic decomposition. In reference [48], the authors introduce the technique of characteristic decomposition for quasi-linear hyperbolic systems, clearly demonstrating the differential relationships among various quantities of the system, and further constructing global solutions for rarefaction wave interactions that occur in high-dimensional Riemann problems, see [24, 35, 42, 43, 44, 46, 47, 49, 50, 52]. We employ their technique to investigate the equations and their corresponding boundary relations. The second technique is the application of symbolic computing software Mathematica. As the incoming and post-shock flows near the limit speeds, due to the significant singularity of various quantities at this time, the calculation is difficult. In this part, we only consider the case where the adiabatic constant is $\gamma=2$. There are two benefits to doing this. First, in this case, the state equation, Bernoulli’s equation, and the Rankine-Hugoniot conditions have polynomial forms, which streamline the calculations. Second, it ensures that various quantities maintain good differentiability near the limit speed. Utilizing computational software Mathematica under these conditions, we are well-equipped to derive the desired computation. We believe that the methods discussed herein can be further extended to yield more general results.
@@
The rest of this paper is organized as follows. In Section 2, we first decompose both the system and boundary conditions by making use of the characteristics and reformulate the problem into one in which $\rho$, $(u,v)$ and their derivatives are expressed by $c$ and the directive derivation along the characteristic, that is ¯±$\bar{\partial}^{\pm}c$. In Section 3, the limit solution is introduced for the case $\underline{u}=\bar{q}$. A particular curve was specified by analysing the jump conditions, the Rankine-Hugoniot, that the shock satisfies. Estimates are obtained here for the states curve $(u,v)$ that could be connected to a given hypersonic flow by a shock by analysing the continuity of a family of parametric curves with parameter relating to the Mach number of the incoming flow. In Section 4.1, we present some a priori estimates for the solution, whose states locate around the limit solution. The proof of Theorem 1.1 is given in Section 4. We establish the a priori estimates for the flow with boundary of Case A in this section and get the global existence of solution. The asymptotic behaviour of the solution is analysed in Section 4.3. Furthermore, we analyse the higher order derivatives of the solution and obtain their limit as the Mach number of the incoming flow goes to infinity in Section 4.4. Finally, for Case B, the proof of Theorem 1.2 is given in Section 5. In Section 5.1, we present the delicate states estimates around the shock polar. Then, under the convex assumption, we show the sign estimates of the solution and prove the global existence. The shock is proved extinguished at infinity for $f_{\infty}=0$.
@@ 2.1. Preliminary
In the following discussion, sometimes for the sake of convenience, we list both $\vec{U}$ and $\epsilon$, but in fact, they share a constraint condition (2.1.2).
@@ 2.2. Characteristic decomposition
Next, we intend to derive the relation of ¯±$\bar{\partial}^{\pm}c$ on the shock $y=\phi(x)$ by (2.1.2). By (1.0.16), define the angle of shock $s$ as follows
@@
$g(\vec{U},\epsilon)$ is continuous with respect to both $\vec{U}$ and $\epsilon$ under the constrain (2.1.2). As previously, we use the notation $g=g(\vec{U},\epsilon)$ in the sequel without confusion. Therefore, Problem 2 induces the derivatives relations (2.2.9), (2.3.2) and (2.4.8).
@@ 3. A limit case
In this section, we introduce the limit solution for the case where $\underline{u}=\bar{q}$ ($\underline{\rho}\underline{u}=\epsilon=0$) and present some useful estimates. This solution is first introduced in [33, 34] to consider the hypersonic flow past a bend obstacle.
@@ 3.1. The limit solution
On the $(u,v)$ plane, it is a circle passing through the origin and centered at $(\frac{\bar{q}}{2},0)$. For such a case, we have many simplifications. First, for the shock $y=\phi(x)$, since $\underline{\rho}=\rho(\bar{q})=\bar{\rho}=0$, it yields
@@
We define a nonzero flow field $(u_{s},v_{s})$ along $\mathsf{W}$ by combining (1.0.14) and (3.1.3) directly as follows
@@
where $S_{o}>0$ is any positive constant. Based on this limit solution, Bae and Xiang establish the global existence of a detached transonic shock for a supersonic flow onto a blunt body in [3].
@@ 3.2. Differential relations for the limit solution
Moreover, the shock polar (2.1.2) becomes the circle (2.4.5), and we have
@@
For the limit solution $\left(u_{s},v_{s}\right)$ in (3.1.5), we can calculate directly that (3.2.4) and (3.2.6) are equivalent to
@@
If $0<\underline{f}<\bar{f}<\sqrt{\frac{2}{\gamma-1}}$, then there exists a positive constant $\delta_{\bar{K}}$, which is depending on $\underline{f}$ and $\bar{f}$, to ensure the following estimate
@@
provided $0<\underline{f}<\bar{f}<\sqrt{\frac{2}{\gamma-1}}$. For any $(u,v)\in\mathcal{C}_{s}^{\delta_{\bar{K}}}$, similar to the proof in Lemma 3.4, we can get the conclusion by the continuity of the function $g(\vec{U},\epsilon)$. ∎